# Removable singularities for Lipschitz fractional caloric functions in time varying domains

### v3

@@ Abstract.
In this paper we study removable singularities for regular $(1,\frac{1}{2s})$-Lipschitz solutions of the $s$-fractional heat equation for $11/2<s<$ 1. To do so, we define a Lipschitz fractional caloric capacity and study its critical dimension and the $L^{2}$-boundedness of a pair of singular integral operators, whose kernels will be the gradient of the fundamental solution of the fractional heat equation and its conjugate.
@@
The author has been supported by PID2020-114167GB-I00 (Mineco, Spain).
@@ 1. Introduction
In the article [MPrT], Mateu, Prat and Tolsa introduce the notion of Lipschitz caloric removability. To define it, objects from parabolic theory in time varying domains are required, such as the parabolic distance, the parabolic BMO space or Lipschitz parabolic functions (see [Ho], [HoL], [NyS]). Inspired by [MPr], we extend these ideas to the fractional caloric setting, giving rise to Lipschitz $s$-parabolic functions as well as a characterization of Lipschitz $s$-caloric removability using a certain capacity. Intuitively, removable sets can be thought as those which “do not matter” when finding solutions of the $s$-heat equation with the previous regularity assumptions. That is, any $s$-parabolic Lipschitz function solving the $s$-heat equation outside the set actually verifies the same equation throughout the entire domain, including the set itself.
Let us introduce basic notation and terminology. Our ambient space will be $\mathbb{R}^{n+1}$ with a generic point denoted by $\overline{x}:=(x,t)\in\mathbb{R}^{n}\times\mathbb{R}$. For each $s\in(0,1]$, the $s$-heat operator is defined as
@@
The fundamental solution $P_{s}(x,t)$ of the $\Theta^{s}$-equation is given by the inverse spatial Fourier transform of $e^{-4\pi^{2}t|\xi|^{2s}}$ for $t>0$, and it equals 00 if $t\leq 0$. If $s=1$, the function $P_{1}$ is nothing but the classical heat kernel,
@@
with $\phi_{n,1}(\rho):=e^{-\rho^{2}/4}$. Although $P_{s}$ is not explicit in general, Blumenthal and Getoor [BG, Theorem 2.1] proved that for $s<1$,
@@
It is not hard to prove that, by construction, one has
@@
With all the above tools and taking $s=1$ (so that $(-\Delta)^{s}$ becomes the usual spatial Laplacian), in [MPrT] the authors introduce the so-called Lipschitz caloric capacity to characterize removable sets for solutions of the heat equation satisfying a Lipschitz condition in space and a $\text{Lip}_{1/2}$ condition in time. That is, functions $f:\mathbb{R}^{n+1}\to\mathbb{R}$ satisfying $\|\nabla_{x}f\|_{\infty}\lesssim 1$ and $\|f\|_{\text{Lip}_{1/2,t}}\lesssim 1$. Nevertheless, in [MPrT] functions are such that
$\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}\lesssim 1\qquad\text{and}\qquad\|\partial_{t}^{1/2}f\|_{\ast,p_{1}}\lesssim 1.$
Using [Ho, Lemma 1] and [HoL, Theorem 7.4], it becomes clear that the previous bounds imply the desired $(1,1/2)$-Lipschitz property, so that the above normalization becomes, in fact, more general. In this paper we extend this to the $s$-caloric setting by introducing a corresponding capacity $\Gamma_{\Theta^{s}}$, for $(1,\frac{1}{2s})$-Lipschitz solutions $f$ of the $\Theta^{s}$-equation, $11/2<s<$ 1, i.e.
$\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}\lesssim 1\qquad\text{and}\qquad\|f\|_{\text{Lip}_{\frac{1}{2s},t}}\lesssim 1.$
We begin our analysis in §2 by noting that the results of Hofmann and Lewis [HoL] can be adapted to the fractional caloric case. We prove:
@@
The above theorem allows us to define $\Gamma_{\Theta^{s}}$ as in [MPrT]. For instance, given a compact set $E\subset\mathbb{R}^{n+1}$, $\Gamma_{\Theta^{s}}$ is
@@
In §3 we localize the potentials associated with the kernels $\nabla_{x}P_{s}$ and $\partial_{t}^{\frac{1}{2s}}P_{s}$ in order to prove (in §4) the equivalence between the removability of compact sets for $(1,\frac{1}{2s})$-Lipschitz solutions of the $\Theta^{s}$-equation and the nullity of $\Gamma_{\Theta^{s}}$. More precisely, a compact set $E\subset\mathbb{R}^{n+1}$ is said to be Lipschitz $s$-caloric removable if for any open subset $\Omega\subset\mathbb{R}^{n+1}$, any function $f:\mathbb{R}^{n+1}\to\mathbb{R}$ with $\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}<\infty$ and $\|\partial_{t}^{\frac{1}{2s}}f\|_{\ast,p_{s}}<\infty$ satisfying the $\Theta^{s}$-equation in $\Omega\setminus{E}$, also satisfies the previous equation in $\Omega$. The main theorem of this article reads as follows:
@@
Moreover, we establish that in $\mathbb{R}^{n+1}$ the critical $s$-parabolic Hausdorff dimension of the previous capacity is $n+1$ and provide an example, for each $s\in(1/2,1)$, of a subset with positive $\pazocal{H}_{p_{s}}^{n+1}$-measure and null capacity. Such set will be a generalization of the Cantor set defined in [MPrT, §6] that, as $s$ approaches the value $1/2$, becomes progressively more and more dense in the unit cube.
@@
Finally, in §5, motivated by a question posed by X. Tolsa, we revisit the proper parabolic case $s=1$ and define a new capacity, different to that of [MPrT], now working with solutions of the heat equation satisfying
@@
In §2, §3 and §4 we fix $s\in(1/2,1)$ The forthcoming results for the $1$-parabolic case are already covered in [MPrT].
@@
$\big{(}I_{p_{s}}f\big{)}^{\wedge}(\overline{\xi}):=\|\overline{\xi}\|_{p_{s}}^{-1}\widehat{f}(\overline{\xi}),$
@@
$\big{(}D_{p_{s}}f\big{)}^{\wedge}(\overline{\xi}):=\|\overline{\xi}\|_{p_{s}}\widehat{f}(\overline{\xi}).$
@@
$(R_{j,s})^{\wedge}:=\frac{\xi_{j}}{\|\overline{\xi}\|_{p_{s}}},\qquad j=1,\ldots,n,$
are the $s$-parabolic Riesz transforms. As observed in the comments that follow [HoL, Equation (2.10)], operators $R_{j,s}$ satisfy analogous properties to those of usual Riesz transforms. Indeed, observe that their symbols are antisymmetric and $s$-parabolically homogeneous of degree 00. Thus, $R_{j,s}$ are defined via convolution (in a principal value sense) against odd kernels, $s$-parabolically homogeneous of degree $-n-2s$, which are bounded in $L^{p}$ and $\text{BMO}_{p_{s}}$ (see [Pe, Remark 1.3]). We have also set $\overline{\xi}:=(\xi,\tau)$ and defined
$(R_{n+1,s})^{\wedge}(\overline{\xi}):=\frac{\tau}{\|\overline{\xi}\|_{p_{s}}^{2s}},\qquad(D_{n+1,s})^{\wedge}(\overline{\xi}):=\frac{\tau}{\|\overline{\xi}\|_{p_{s}}^{2s-1}}.$
@@
$(D_{n+1,s}f)^{\wedge}(\overline{\xi})=\Big{[}|\tau|^{\frac{1}{2s}}m(\overline{\xi})\Big{]}\widehat{f},\qquad\text{where }\quad m(\overline{\xi}):=\frac{\tau}{|\tau|^{\frac{1}{2s}}\|\overline{\xi}\|_{p_{s}}^{2s-1}}.$
Now, since $m$ is not smooth in $\mathbb{R}^{n+1}\setminus{\{0\}}$, let us consider an even function $\phi\in\pazocal{C}^{\infty}(\mathbb{R})$, which equals 1 on $(2/K,\infty)$ and is supported on $(-\infty,-1/K)\cup(1/K,\infty)$, for some $K\geq 2$. We also choose $\phi$ so that $\|\partial^{l}\phi\|_{\infty}\lesssim K^{l}$, for $50\leq l\leq n+$ 5. Let us introduce the following auxiliary functions,
@@
This way, we rewrite $(D_{n+1,s}f)^{\wedge}$ as follows:
$(D_{n+1,s}f)^{\wedge}(\overline{\xi})$ $=\bigg{[}m^{+}(\overline{\xi})|\tau|^{\frac{1}{2s}}+m(\overline{\xi})\big{(}1-\phi\big{)}\bigg{(}\frac{\tau}{|\xi|^{2s}}\bigg{)}|\tau|^{\frac{1}{2s}}\bigg{]}\widehat{f}(\overline{\xi})$
@@
Regarding $m^{+}$, observe that it is an odd multiplier, smooth in $\mathbb{R}^{n+1}\setminus{\{0\}}$ and $s$-parabolically homogeneous of degree 0. Therefore, by [Gr, Proposition 2.4.7] (which admits a direct adaptation to the $s$-parabolic case), $m^{+}$ is associated to a convolution kernel $L^{+}$, odd, $s$-parabolically homogeneous of degree $-n-2s$ and bounded on $L^{p}$ and $\text{BMO}_{p_{s}}$ (see again [Pe, Remark 1.3]).
@@
$\bigg{\rvert}\int_{\mathbb{R}}\frac{\sigma(\xi,\tau)-\sigma(\xi,r)}{|t-r|^{1+\frac{1}{4s}}}\mathop{}\!\mathrm{d}r\bigg{\rvert}\begin{dcases}=0,&\text{if }\,|\xi|\geq C2^{-i},\\ \lesssim\frac{2^{in}}{|\tau|^{1+\frac{1}{4s}}},&\text{if }\,|\tau|>\frac{4|\xi|^{2s}}{K},\\ \lesssim\frac{2^{i(n+2s-1)}}{|\tau|^{\frac{3}{4s}}},&\text{if }\,|\tau|\leq\frac{8|\xi|^{2s}}{K}\leq C^{\prime}2^{-2si},\\ \end{dcases}$
@@ 3. Localization of potentials
|II$(\overline{x})|| I$ $\leq\int_{2Q}|\nabla_{x}P_{s}(\overline{x}-\overline{y})||g_{1}(\overline{y})|\mathop{}\!\mathrm{d}\overline{y}+\int_{\mathbb{R}^{n+1}\setminus{2Q}}|\nabla_{x}P_{s}(\overline{x}-\overline{y})||g_{1}(\overline{y})|\mathop{}\!\mathrm{d}\overline{y}$
@@
where we have chosen $10<\varepsilon<2s-$ 1. On the other hand, if $\overline{x}\notin 4Q$,
@@
(3.3) $\hskip 28.45274pt+P_{s}\ast I_{s}(\varphi,P_{s}\ast T)(\overline{x})-P_{s}\ast I_{s}(\varphi,P_{s}\ast T)(\widetilde{x})$
(3.4) $\hskip 28.45274pt-P_{s}\ast\big{(}(P_{s}\ast T-c)\Theta^{s}\varphi\big{)}(\overline{x})+P_{s}\ast\big{(}(P_{s}\ast T-c)\Theta^{s}\varphi\big{)}(\widetilde{x}).$
@@
[…]
@@
$\leq\ell(Q)^{2s}\sum_{j=1}^{\infty}\int_{(\mathbb{R}^{n}\times I_{Q})\cap A_{j}}\frac{\mathop{}\!\mathrm{d}y}{|x-y|^{n+1}}\lesssim\ell(Q)^{2s}\sum_{j=1}^{\infty}\frac{(2^{j}\ell(Q))^{n}}{(2^{j}\ell(Q))^{n+1}}=\ell(Q)^{2s-1}.$
@@
$\bigg{\rvert}\frac{1}{|r-t|^{1+\frac{1}{2s}}}-\frac{1}{|r-u|^{1+\frac{1}{2s}}}\bigg{\rvert}|\lesssim\frac{|t-u|}{|r-t|^{2+\frac{1}{2s}}}.$
@@
So once (3.8) is proved, case i) is done. To prove it, we split $\mathbb{R}^{n+1}\setminus{4Q}$ into $s$-parabolic annuli $A_{k}:=2^{k+1}Q\setminus 2^{k}Q$ and consider $\pazocal{C}^{\infty}$ functions $\widetilde{\chi}_{k}$ such that $\chi_{A_{k}}\leq\widetilde{\chi}_{k}\leq\chi_{\frac{11}{10}A_{k}}$, as well as $\|\nabla_{x}\widetilde{\chi}_{k}\|_{\infty}\lesssim(2^{k}\ell(Q))^{-1}$, $\|\partial_{t}\widetilde{\chi}_{k}\|_{\infty}\lesssim(2^{k}\ell(Q))^{-2s}$ and
$\sum_{k\geq 2}\widetilde{\chi}_{k}=1,\qquad\text{on }\,\mathbb{R}^{n+1}\setminus{4Q}.$
@@
In order to prove 2 we apply an $s$-parabolic version of Frostman’s lemma. Let us name $d:=\text{{dim}}_{\pazocal{H}_{p_{s}}}(E)$ and assume $0<\pazocal{H}_{p_{s}}^{d}(E)<\infty$ without loss of generality. Indeed, if $\pazocal{H}_{p_{s}}^{d}(E)=\infty$, apply an $s$-parabolic version of [F, Theorem 4.10] to construct a compact set $\widetilde{E}\subset\mathbb{R}^{n+1}$ with $\widetilde{E}\subset E$ and $0<\pazocal{H}_{p_{s}}^{d}(\widetilde{E})<\infty$. On the other hand, if $\pazocal{H}_{p_{s}}^{d}(E)=0$ apply the same reasoning with $d^{\prime}:=(d+n+1)/2$. In any case, by Frostman’s lemma we shall then consider a non trivial positive measure $\mu$ with $\text{supp}(\mu)\subset E$ satisfying $\mu(B(\overline{x},r))\leq r^{d}$, for all $\overline{x}\in\mathbb{R}^{n+1}$ and all $r>0$. Observe that if we prove
@@
[…]
@@
$\pazocal{P}^{s}_{\mu,\varepsilon}f(\overline{x}):=\int_{|\overline{x}-\overline{y}|>\varepsilon}P_{s}(\overline{x}-\overline{y})f(\overline{y})\text{d}\mu(\overline{y}),\hskip 14.22636pt\overline{x}\in\mathbb{R}^{n+1},\;\varepsilon>0.$
@@
Let us fix $10<\varepsilon<$ 1 and $\overline{x}=(x,t)\in\mathbb{R}^{n+1}$ and notice that for some $10<\delta<2s-$ 1,
@@
$\leq\int_{B(\overline{x},\varepsilon)}\bigg{(}\int_{B_{1}(y,2\varepsilon)}\frac{\mathop{}\!\mathrm{d}z}{|z-y|^{n-\delta}}\int_{-2^{2s}\varepsilon^{2s}}^{2^{2s}\varepsilon^{2s}}\frac{du}{u^{\frac{1+\delta}{2s}}}\bigg{)}\text{d}|\mu|(\overline{y})\lesssim\varepsilon^{\delta}\varepsilon^{2s-1-\delta}|\mu|(B(\overline{x},\varepsilon))\lesssim\varepsilon^{n+2s}.$
@@
$|\displaystyle\big{\rvert}\pazocal{P}_{\varepsilon}^{s}\mu(\overline{x})-\pazocal{P}^{s}\mu(\overline{z})\big{\rvert}$$\leq\big{\rvert}\pazocal{P}_{\varepsilon}^{s}\mu(\overline{x})-\pazocal{P}_{\varepsilon}^{s}\mu(\overline{z})\big{\rvert}+C_{2}$
@@
[…]

### v4

@@ Abstract.
In this paper we study removable singularities for regular $(1,\frac{1}{2s})$-Lipschitz solutions of the $s$-fractional heat equation for $1/2<s<1$. To do so, we define a Lipschitz fractional caloric capacity and study its critical dimension and the $L^{2}$-boundedness of a pair of singular integral operators, whose kernels will be the gradient of the fundamental solution of the fractional heat equation and its conjugate.
@@
ORCID: 0000-0002-2207-5981. The author has been supported by PID2020-114167GB-I00 (Mineco, Spain).
@@ 1. Introduction
In the article [MPrT], Mateu, Prat and Tolsa introduce the notion of Lipschitz caloric removability. To define it, objects from parabolic theory in time varying domains are required, such as the parabolic distance, the parabolic BMO space, or Lipschitz parabolic functions (see [Ho], [HoL], [NyS]). Inspired by [MPr], we extend these ideas to the fractional caloric setting, giving rise to Lipschitz $s$-parabolic functions, as well as a characterization of Lipschitz $s$-caloric removability using a certain capacity. Intuitively, removable sets can be thought of as those that do not affect the search for solutions of the $s$-heat equation with the previous regularity assumptions. That is, any $s$-parabolic Lipschitz function solving the $s$-heat equation outside the set actually verifies the same equation throughout the entire domain, including the set itself.
We introduce basic notation and terminology. The ambient space will be $\mathbb{R}^{n+1}$ with a generic point denoted by $\overline{x}:=(x,t)\in\mathbb{R}^{n}\times\mathbb{R}$. For each $s\in(0,1]$, the $s$-heat operator is defined as
@@
The fundamental solution $P_{s}(x,t)$ of the $\Theta^{s}$-equation is given by the inverse spatial Fourier transform of $e^{-4\pi^{2}t|\xi|^{2s}}$ for $t>0$, and it equals 0 if $t\leq 0$. If $s=1$, the function $P_{1}$ is nothing but the classical heat kernel,
@@
with $\phi_{n,1}(\rho):=e^{-\rho^{2}/4}$. Blumenthal and Getoor [BG, Theorem 2.1] proved that for $s<1$,
@@
It follows from the construction that
@@
For $s=1$ (the classical caloric setting) a function $f$ is said to be $(1,1/2)$-Lipschitz regular if it is such that
(1.1) $\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}<\infty,\hskip 28.45274pt\|\partial_{t}^{1/2}f\|_{\ast,p_{1}}<\infty.$
As shown by Hofmann and Lewis [Ho, Lemma 1], [HoL, Thm. 7.4], these functions satisfy
$\|f\|_{\text{Lip}_{1/2},t}:=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{n}\\ t,u\in\mathbb{R},t\neq u\end{subarray}}\frac{|f(x,t)-f(x,u)|}{|t-u|^{1/2}}\lesssim\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}+\|\partial_{t}^{1/2}f\|_{\ast,p_{1}}.$
Thus a $(1,1/2)$-Lipschitz function is Lipschitz in the spatial variables and 1/2-Lipschitz in time. In [MPrT] the authors introduce the so-called Lipschitz caloric capacity to characterize removable sets for solutions of the heat equation satisfying the above $(1,1/2)$--Lipschitz condition
The main goal of this paper is to extend the previous notions and results to the fractional caloric setting. Namely, we will introduce a corresponding capacity $\Gamma_{\Theta^{s}}$ for $(1,\frac{1}{2s})$-Lipschitz solutions $f$ of the $\Theta^{s}$-equation, $1/2<s<1$, i.e. those satisfying
$\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}\lesssim 1\qquad\text{and}\qquad\|f\|_{\text{Lip}_{\frac{1}{2s},t}}:=\sup_{\begin{subarray}{c}x\in\mathbb{R}^{n}\\ t,u\in\mathbb{R},t\neq u\end{subarray}}\frac{|f(x,t)-f(x,u)|}{|t-u|^{\frac{1}{2s}}}\lesssim 1.$
We begin our analysis in §2 by noting that the results of Hofmann and Lewis [HoL] can be adapted to the fractional caloric case:
@@
The proof of the above result follows the arguments of [HoL] and adapts them to the fractional parabolic setting. For instance, one needs to work with $s$-parabolic Riesz transforms instead of the usual ones from the Euclidean case. Moreover, we need to follow the dependence on $s$ when establishing the above estimates and, in fact, in Remark 2.1 we observe that the arguments break down when $s=1/2$, which is consistent with the theory.
This theorem allows us to define the capacity $\Gamma_{\Theta^{s}}$ in analogy with the classical case [MPrT]. For instance, given a compact set $E\subset\mathbb{R}^{n+1}$, $\Gamma_{\Theta^{s}}$ is
@@
In §3 we localize the potentials associated with the kernels $\nabla_{x}P_{s}$ and $\partial_{t}^{\frac{1}{2s}}P_{s}$. Although our methods of proof and statement of results will be analogous to those of [MPrT] for the classical heat case, we will need to adapt many techniques and generalize some of the methods used in the previous reference. Indeed, we need to take into account that the operator $\Theta^{s}$ for $s<1$ is no longer local, so many properties will no longer hold. For example, if $s=1$ one can exploit the useful relation
$\Delta(fg)=g\Delta f+f\Delta g+2\nabla f\cdot\nabla g.$
However, if $s<1$ one has to work with expressions of the form
$(-\Delta)^{s}(fg)=f(-\Delta)^{s}g+g(-\Delta)^{s}f-I_{s}(f,g),$
where $I_{s}$ is a crossed-terms integral defined as
In §4 we use the localization of potentials to prove the equivalence between the removability of compact sets for $(1,\frac{1}{2s})$-Lipschitz solutions of the $\Theta^{s}$-equation and the nullity of $\Gamma_{\Theta^{s}}$. More precisely, a compact set $E\subset\mathbb{R}^{n+1}$ is said to be Lipschitz $s$-caloric removable if for any open subset $\Omega\subset\mathbb{R}^{n+1}$, any function $f:\mathbb{R}^{n+1}\to\mathbb{R}$ with $\|\nabla_{x}f\|_{L^{\infty}(\mathbb{R}^{n+1})}<\infty$ and $\|\partial_{t}^{\frac{1}{2s}}f\|_{\ast,p_{s}}<\infty$ satisfying the $\Theta^{s}$-equation in $\Omega\setminus{E}$, also satisfies the previous equation in $\Omega$. The main theorem of this article reads as follows:
@@
Moreover, we establish that in $\mathbb{R}^{n+1}$ the critical $s$-parabolic Hausdorff dimension of the previous capacity is $n+1$ and provide an example, for each $s\in(1/2,1)$, of a subset with positive $\pazocal{H}_{p_{s}}^{n+1}$-measure and null capacity. This already shows that the capacity $\Gamma_{\Theta^{s}}$ has to be fundamentally different to those studied by the author in [HeMPr], since it cannot be comparable to the $s$-parabolic Hausdorff content $\pazocal{H}_{\infty,p_{s}}^{n+1}$. Such set will be a generalization of the Cantor set defined in [MPrT, §6] whose construction will heavily depend on the parameter $s$. In fact, as $s$ approaches the value $1/2$, our Cantor set will become progressively more and more dense (even for the first iteration of its construction) in the unit cube.
@@
Finally, in §5, motivated by a question posed by X. Tolsa, we revisit the classical parabolic case $s=1$ and define a new capacity, different to that of [MPrT], now working with solutions of the heat equation satisfying
@@
In §2, §3 and §4 we fix $s\in(1/2,1)$ . The forthcoming results for the $1$-parabolic case are already covered in [MPrT].
@@
$\widehat{I_{p_{s}}f}(\overline{\xi}):=\|\overline{\xi}\|_{p_{s}}^{-1}\widehat{f}(\overline{\xi}),$
@@
$\widehat{D_{p_{s}}f}(\overline{\xi}):=\|\overline{\xi}\|_{p_{s}}\widehat{f}(\overline{\xi}).$
@@
$\widehat{R_{j,s}}:=\frac{\xi_{j}}{\|\overline{\xi}\|_{p_{s}}},\qquad j=1,\ldots,n,$
are the $s$-parabolic Riesz transforms. As observed in the comments that follow [HoL, Equation (2.10)], operators $R_{j,s}$ satisfy analogous properties to those of usual Riesz transforms. Indeed, observe that their symbols are antisymmetric and $s$-parabolically homogeneous of degree 0. Thus, $R_{j,s}$ are defined via convolution (in a principal value sense) against odd kernels, $s$-parabolically homogeneous of degree n2s-n-2s, which are bounded in $L^{p}$ and $\text{BMO}_{p_{s}}$ (see [Pe, Remark 1.3]). We have also set $\overline{\xi}:=(\xi,\tau)$ and defined
$\widehat{R_{n+1,s}}(\overline{\xi}):=\frac{\tau}{\|\overline{\xi}\|_{p_{s}}^{2s}},\qquad\widehat{D_{n+1,s}}(\overline{\xi}):=\frac{\tau}{\|\overline{\xi}\|_{p_{s}}^{2s-1}}.$
@@
$\widehat{D_{n+1,s}}f(\overline{\xi})=\Big[|\tau|^{\frac{1}{2s}}m(\overline{\xi})\Big]\widehat{f},\qquad\text{where }\quad m(\overline{\xi}):=\frac{\tau}{|\tau|^{\frac{1}{2s}}\|\overline{\xi}\|_{p_{s}}^{2s-1}}.$
Now, since $m$ is not smooth in $\mathbb{R}^{n+1}\setminus{\{0\}}$, let us consider an even function $\phi\in\pazocal{C}^{\infty}(\mathbb{R})$, which equals 1 on $(2/K,\infty)$ and is supported on $(-\infty,-1/K)\cup(1/K,\infty)$, for some $K\geq 2$. We also choose $\phi$ so that $\|\partial^{l}\phi\|_{\infty}\lesssim K^{l}$, for $0\leq l\leq n+5$. Let us introduce the following auxiliary functions,
@@
This way, we rewrite $\widehat{D_{n+1,s}f}$ as follows:
$\widehat{D_{n+1,s}f}(\overline{\xi})$ $=\bigg[m^{+}(\overline{\xi})|\tau|^{\frac{1}{2s}}+m(\overline{\xi})\big(1-\phi\big)\bigg(\frac{\tau}{|\xi|^{2s}}\bigg)|\tau|^{\frac{1}{2s}}\bigg]\widehat{f}(\overline{\xi})$
@@
Regarding $m^{+}$, observe that it is an odd multiplier, smooth in $\mathbb{R}^{n+1}\setminus{\{0\}}$ and $s$-parabolically homogeneous of degree 0. Therefore, by [Gr, Proposition 2.4.7] (which admits a direct adaptation to the $s$-parabolic case), $m^{+}$ is associated to a convolution kernel $L^{+}$, odd, $s$-parabolically homogeneous of degree n2s-n-2s and bounded on $L^{p}$ and $\text{BMO}_{p_{s}}$ (see again [Pe, Remark 1.3]).
@@
[…]
@@ 3. Localization of potentials
$|\text{II}(\overline{x})|$ $\leq\int_{2Q}|\nabla_{x}P_{s}(\overline{x}-\overline{y})||g_{1}(\overline{y})|\mathop{}\!\mathrm{d}\overline{y}+\int_{\mathbb{R}^{n+1}\setminus{2Q}}|\nabla_{x}P_{s}(\overline{x}-\overline{y})||g_{1}(\overline{y})|\mathop{}\!\mathrm{d}\overline{y}$
@@
where we have chosen $0<\varepsilon<2s-1$. On the other hand, if $\overline{x}\notin 4Q$,
@@
[…]
@@
In order to estimate the remaining differences (3.3) and (3.4) we name $g_{1}:=(P_{s}\ast T-c)\Theta^{s}\varphi$ and $g_{2}:=I_{s}(\varphi,P_{s}\ast T)$. Such functions have already appeared in the proof of Lemma 3.2 and satisfy $\text{supp}(g_{j})\subset\mathbb{R}^{n}\times I_{Q}$ and $\|g_{j}\|_{\infty}\lesssim\ell(Q)^{-2s+1}$, for $j=1,2$. By a direct application of Lemma 3.3 we deduce that both differences are bounded by $|r-t|^{\frac{1}{2s}}$ up to a constant, and we are done. ∎
@@
$\hskip-21.33955pt\leq\ell(Q)^{2s}\sum_{j=1}^{\infty}\int_{(\mathbb{R}^{n}\times I_{Q})\cap A_{j}}\frac{\mathop{}\!\mathrm{d}y}{|x-y|^{n+1}}\lesssim\ell(Q)^{2s}\sum_{j=1}^{\infty}\frac{(2^{j}\ell(Q))^{n}}{(2^{j}\ell(Q))^{n+1}}=\ell(Q)^{2s-1}.$
@@
$\bigg\rvert\frac{1}{|r-t|^{1+\frac{1}{2s}}}-\frac{1}{|r-u|^{1+\frac{1}{2s}}}\bigg\rvert\lesssim\frac{|t-u|}{|r-t|^{2+\frac{1}{2s}}}.$
@@
So once (3.8) is proved, case i) is done. To prove it, we split $\mathbb{R}^{n+1}\setminus{4Q}$ into $s$-parabolic annuli $A_{k}:=2^{k+1}Q\setminus 2^{k}Q$ and consider $\pazocal{C}^{\infty}$ functions $\widetilde{\chi}_{k}$ such that $\chi_{A_{k}}\leq\widetilde{\chi}_{k}\leq\chi_{\frac{11}{10}A_{k}}$, as well as $\|\nabla_{x}\widetilde{\chi}_{k}\|_{\infty}\lesssim(2^{k}\ell(Q))^{-1}$, $\|\partial_{t}\widetilde{\chi}_{k}\|_{\infty}\lesssim(2^{k}\ell(Q))^{-2s}$ and $\sum_{k\geq 2}\widetilde{\chi}_{k}=1$ on $\mathbb{R}^{n+1}\setminus{4Q}$.
@@
[…]
@@
$\pazocal{P}_{\mu}^{s}f(\overline{x}):=\int_{\mathbb{R}^{n+1}}\nabla_{x}P_{s}(\overline{x}-\overline{y})f(\overline{y})\text{d}\mu(\overline{y}),\hskip 14.22636pt\overline{x}\notin\text{supp}(\mu).$
@@
$\pazocal{P}^{s}_{\mu,\varepsilon}f(\overline{x}):=\int_{|\overline{x}-\overline{y}|>\varepsilon}\nabla_{x}P_{s}(\overline{x}-\overline{y})f(\overline{y})\text{d}\mu(\overline{y}),\hskip 14.22636pt\overline{x}\in\mathbb{R}^{n+1},\;\varepsilon>0.$
@@
Let us fix $0<\varepsilon<1$ and $\overline{x}=(x,t)\in\mathbb{R}^{n+1}$ and notice that for some $0<\delta<2s-1$,
@@
$\hskip-78.24507pt\leq\int_{B(\overline{x},\varepsilon)}\bigg(\int_{B_{1}(y,2\varepsilon)}\frac{\mathop{}\!\mathrm{d}z}{|z-y|^{n-\delta}}\int_{-2^{2s}\varepsilon^{2s}}^{2^{2s}\varepsilon^{2s}}\frac{du}{u^{\frac{1+\delta}{2s}}}\bigg)\text{d}|\mu|(\overline{y})\lesssim\varepsilon^{\delta}\varepsilon^{2s-1-\delta}|\mu|(B(\overline{x},\varepsilon))\lesssim\varepsilon^{n+2s}.$
@@
[…]
@@
$A_{j}:=\big\{\overline{y}\,:\,2^{j-1}\varepsilon\leq|\overline{x}-\overline{y}|_{p_{s}}\leq 2^{j}\varepsilon\big\},\quad j\geq 0,$